# TARS [Uncaptioned image]: MinMax Token-Adaptive Preference Strategy for Hallucination Reduction in MLLMs

### v2

@@ Abstract
Multimodal large language models (MLLMs) enable vision-language reasoning, yet often generate plausible outputs that are factually incorrect or visually ungrounded, thereby compromising their reliability. Direct preference optimization (DPO) is a common strategy for correcting hallucinations by aligning model outputs with human preferences. Existing DPO strategies typically treat hallucination-related preferences as fixed targets, relying on static supervision signals during training. This approach tends to overfit to superficial linguistic cues in preference data, leading to distributional rigidity and spurious correlations that impair grounding in causally relevant visual information. To overcome this limitation, we propose TARS, a token-adaptive preference strategy that reformulates DPO as a min-max optimization problem. TARS maximizes token-level distributional shifts under semantic constraints to simulate alignment uncertainty, and simultaneously minimizes the expected preference loss under these controlled perturbations. This joint objective preserves causal grounding while mitigating overfitting to preference patterns, thereby reducing hallucinations in multimodal reasoning. We evaluate TARS on multiple hallucination benchmarks and find consistently strong performance. Using only 4.8k preference samples and no expert feedback, TARS reduces hallucination rates from 26.4% to 13.2% and decreases cognition value from 2.5 to 0.4. It outperforms standard DPO and matches GPT-4o on several key metrics.
{strip}
[Uncaptioned image] [Uncaptioned image]
Figure 1: Left: We present TARS, a token-adaptive preference strategy for mitigating hallucinations in MLLMs. TARS reformulates direct preference optimization (DPO) as a min-max objective that (1) minimizes behavioral misalignment via preference feedback and (2) maximizes adaptability through perturbations of visual-agnostic tokens. Right: Evaluation on LLaVA-v1.5-13B with preference optimization (PO) (Liu et al. 2023b) and industrial MLLMs under the AMBER benchmark (Wang et al. 2023) shows that TARS surpasses PO baselines and matches GPT-4o (Hurst et al. 2024) in hallucination suppression.
@@ 1 Introduction
Large language models (LLMs) demonstrate strong reasoning capabilities across a broad range of language tasks (Gandhi et al. 2023; Chen et al. 2023; Wang et al. 2025). Building on this foundation, multimodal large language models (MLLMs) integrate visual inputs to enable grounded understanding and vision-language reasoning (Tong et al. 2024; Huang et al. 2023; Driess et al. 2025). Although this integration broadens their applicability in various tasks (Jiang et al. 2024b; Shao et al. 2025), it also introduces key challenges, among which hallucinations are particularly prominent (Kim et al. 2024; Jiang et al. 2024a; Huang et al. 2024). Hallucinations in MLLMs refer to outputs that may appear plausible yet are factually incorrect or lack grounding in the visual context (Sarkar et al. 2025b; Gunjal, Yin, and Bas 2024). Addressing these failures is essential for improving the reliability, safety, and practical applicability of MLLMs in real-world applications.
@@
Modern MLLMs are typically developed through a two-stage training pipeline that includes a knowledge-intensive pretraining phase (Dai et al. 2023; Bao et al. 2022; Zhang et al. 2024), followed by instruction tuning (Liu et al. 2024a; Wang et al. 2024b; Liu et al. 2023a). These stages endow the model with broad world knowledge and the ability to follow instructions in natural language (Bao et al. 2022; Liu et al. 2023b). Despite these capabilities, hallucinations often stem not from knowledge deficits, but from behavioral biases acquired during training that lead the model to generate plausible yet ungrounded outputs (Oh et al. 2024; Chen et al. 2025). To address these failures, preference optimization (PO) has become a prominent strategy for reducing hallucinations by aligning model outputs with human expectations (Schulman et al. 2017; Achiam et al. 2017). PO fine-tunes models using ranked response pairs derived from either human feedback (Sun et al. 2024; Ouyang et al. 2022) or AI-generated preferences (Yu et al. 2025; Sharma et al. 2024), providing direct supervision to reinforce grounded and faithful responses. Such methods have proven effective in mitigating hallucinations across diverse tasks.
Direct preference optimization (DPO) (Rafailov et al. 2023) has become a widely adopted method for hallucination reduction (Fu et al. 2025; Yang et al. 2025). While generally effective, current DPO methods rely heavily on preference data, which can cause models to overfit to shallow textual cues, such as high-frequency phrases or repetitive patterns in the training set (Huang et al. 2024; Liu, Zheng, and Chen 2024). Prior studies have observed that this overfitting leads MLLMs to generate responses that appear plausible but lack visual grounding, as illustrated in Figure 2(c). In our analysis (Figure 2(d)), we further find that DPO-trained models often assign high preference to outputs containing spurious correlation tokens, including prepositions or frequently mentioned objects, even when these elements are not visually grounded (Xie et al. 2024; Wang et al. 2024a). These observations reveal a core limitation of DPO: its reliance on static preference signals hinders generalization under shifting visual–textual contexts, leading to brittle cross-modal alignment and increased vulnerability to hallucination (Setlur et al. 2024; Fu et al. 2025). This rigidity also prevents the model from adapting to local semantic discrepancies, weakening its ability to ground responses in causally relevant visual information.
@@
To validate the effectiveness of our method, we evaluate TARS on LLaVA-v1.5 (Liu et al. 2023b) at 7B and 13B scales, comparing it against leading preference optimization approaches. Across a comprehensive suite of hallucination benchmarks spanning generative and discriminative tasks, TARS-enhanced LLaVA-v1.5 achieves consistently strong performance and matches GPT-4o (Hurst et al. 2024) in several settings. These results underscore the effectiveness of token-adaptive preference optimization in reducing hallucinations and advancing the trustworthiness of MLLMs.
@@ 2 Preliminaries
Multimodal Large Language Models. MLLMs extend large language models (LLMs) by incorporating visual inputs alongside textual prompts (Zhang et al. 2024). Formally, given an image $x$ and a prompt $q$, the model generates a textual response $y=(y_{1},\dots,y_{l})$ in an autoregressive manner (Liu et al. 2023b):
@@
where $\pi_{\theta}$ denotes the model’s conditional generation policy parameterized by $\theta$. Given a multimodal input of a textual input $q$ and a visual input $x$, the model tokenizes them into discrete sequences: textual tokens $q=\{q_{1},\dots,q_{m}\}$ and visual tokens $x=\{x_{1},\dots,x_{n}\}$. These tokens are mapped to embeddings and fused via cross-attention to integrate semantic signals from both modalities. The resulting context is then used by the decoder to autoregressively generate the output sequence (Dou et al. 2022; Yang et al. 2021).
Direct Preference Optimization. Direct preference optimization (DPO) (Rafailov et al. 2023) is an effective approach for aligning model behavior with human preferences. It bypasses explicit reward models by directly optimizing preferences from pairwise comparisons.
Traditional methods such as reinforcement learning with human feedback (RLHF) (Ouyang et al. 2022) and AI feedback (RLAIF) (Yu et al. 2025) rely on training a scalar reward model $r_{\psi}(x,q,y)$ from preference pairs. This reward model is typically trained using the Bradley-Terry formulation (Bradley and Terry 1952):
@@
$\underset{\pi_{\theta}}{\min}\,\mathbb{E}_{(x,q)\sim\mathcal{D},\,y^{*}\sim\pi_{\theta}(y^{*}\mid x,q)}\Big{[}-\Big{(}r_{\psi}(x,q,y^{*})$ (4)
$\qquad-\alpha\cdot\mathbb{D}_{\mathrm{KL}}\left(\pi_{\theta}(y^{*}\mid x,q)\,\middle\|\,\pi_{\text{ref}}(y^{*}\mid x,q)\right)\Big{)}\Big{]},$
where $\alpha$ controls the strength of regularization, which ensures alignment with the learned preferences without drifting too far from the pretrained behavior. Rather than relying on the explicitly trained reward model, DPO (Rafailov et al. 2023) simplifies the learning process by leveraging the insight that the optimal policy can be expressed in closed form using relative log-likelihoods under $\pi_{\theta}$ and $\pi_{\text{ref}}$:
$\underset{\pi_{\theta}}{\min}\,\mathbb{E}_{(x,q,y_{w},y_{r})\sim\mathcal{D}}\Big{[}-\log$ $\sigma\Big{(}\alpha\log\frac{\pi_{\theta}(y_{w}\mid x,q)}{\pi_{\text{ref}}(y_{w}\mid x,q)}$ (5)
$-\alpha\log\frac{\pi_{\theta}(y_{r}\mid x,q)}{\pi_{\text{ref}}(y_{r}\mid x,q)}\Big{)}\Big{]},$
@@ 3 Method
Figure 3: Overview of TARS. TARS reformulates preference optimization as a Min–Max problem: (1) The maximization branch perturbs visual-agnostic tokens to simulate semantically shifted contexts (red dashed box); (2) The minimization branch fine-tunes the model to align with human preferences via the DPO objective (purple dashed box). TARS encourages the model to attend to causally grounded visual signals rather than spurious textual correlations, thereby reducing hallucinations.
@@ 3.2 Maximizing with Token Perturbations
As shown in Equation 5, DPO aligns models with preferred responses via log-likelihood ratios against a reference model. However, we observe that this formulation can encourage overfitting to superficial patterns such as frequent phrases, stylistic tokens, which we find reduce effective alignment with the visual context in multimodal settings (Setlur et al. 2024; Fu et al. 2025).
@@
where $\Phi(\mathcal{A})$ denotes allowable perturbations constrained to $\mathcal{A}(x,q)$, and $\mathrm{Sim}(\varphi(q),q)$ measures token-level deviation. In practice, we approximate $\varphi^{*}(q)$ by applying local token-level transformations:
@@
[…]
@@ 3.3 Spectral Regularization for Token Alignment
While our method introduces token-level perturbations to simulate distribution shifts, the supervision derived from preference pairs $(y_{w},y_{r})$ is static. This discrepancy between adaptive input representations and fixed supervision may encourage the model to learn distribution-specific artifacts, especially under strong alignment constraints (Fu et al. 2025; Chowdhury, Kini, and Natarajan 2024).
In practice, semantic alignment does not require strict token-level correspondence. Enforcing such fine-grained constraints may reintroduce spurious correlations that our min-max strategy aims to mitigate (Zhou et al. 2024; Tian et al. 2025). To address this, we propose frequency-domain alignment, where local token perturbations translate into smooth global variations in the spectral space. This approach ensures semantic consistency between perturbed and original inputs without rigid token-wise matching.
Specifically, we extract hidden states for $(x,\varphi(q),y_{w})$ and compare them to those of $(x,q,y_{w})$ and $(x,q,y_{r})$ using the Fast Fourier Transform (FFT) (Cooley and Tukey 1965). Formally, let $z\in\mathbb{R}^{L\times D}$ denote a sequence of hidden states, where $L$ is the token length. We compute the spectral representation via:
@@
$\mathcal{L}_{\text{freq}}=$ $-\log\sigma\Big{(}\beta\cdot\Big{[}\,\log\frac{\mathcal{F}\left(h_{\theta}(x,\varphi(q),y_{w})\right)}{\mathcal{F}\left(h_{\text{ref}}(x,q,y_{w})\right)}$ (11)
$-\log\frac{\mathcal{F}\left(h_{\theta}(x,\varphi(q),y_{w})\right)}{\mathcal{F}\left(h_{\text{ref}}(x,q,y_{r})\right)}\,\Big{]}\Big{)},$
@@ 3.4 Minimization Objective in TARS
$\mathcal{L}_{\text{TARS}}(x,q,\varphi(q),y_{w},$ $y_{r})=\;\mathcal{L}_{\text{DPO}}(x,\varphi(q),y_{w},y_{r})$ (12)
$+\lambda\cdot\mathcal{L}_{\text{freq}}(x,q,\varphi(q),y_{w},y_{r}),$
@@
Table 1: Comparison across hallucination benchmarks. We evaluate state-of-the-art MLLMs as reference baselines, denoted by $\S$. For algorithms with available checkpoints, results from re-testing are marked with $\dagger$; for those without, we reproduce results using settings from (Fu et al. 2025; Li, Lin, and Pei 2024), denoted by $\ddagger$. All experiments use greedy decoding with temperature set to 0 for consistency and reproducibility. Bold denotes the best performance, and underlined denotes the second-best.

### v3

# TARS [Uncaptioned image]: MinMax Token-Adaptive Preference Strategy for MLLM Hallucination Reduction

@@ Abstract
Multimodal large language models (MLLMs) enable vision-language reasoning, yet often generate plausible outputs that are factually incorrect or visually ungrounded, thereby compromising their reliability. Direct preference optimization (DPO) is a common strategy for correcting hallucinations by aligning model outputs with human preferences. However, existing DPO strategies typically treat hallucination-related preferences as fixed targets, relying on static and potentially biased supervision signals during training. This approach tends to overfit to superficial linguistic cues in preference data, leading to distributional rigidity and spurious correlations that impair grounding in causally relevant visual information. To overcome this limitation, we propose TARS, a token-adaptive preference strategy that reformulates DPO as a min–max optimization problem. TARS maximizes token-level distributional shifts under explicit semantic constraints to simulate alignment uncertainty, and simultaneously minimizes the expected preference loss under these controlled perturbations. This joint objective effectively preserves causal grounding while mitigating overfitting to preference patterns, thereby reducing hallucinations in multimodal reasoning. We evaluate TARS on multiple hallucination benchmarks and find consistently strong and robust performance. Using only 4.8k preference samples and no expert feedback, TARS reduces hallucination rates from 26.4% to 13.2% and decreases cognition value from 2.5 to 0.4, outperforming standard DPO and matching GPT-4o on several key metrics.
Refer to caption Refer to caption
Figure 1: Left: We present TARS, a token-adaptive preference strategy for mitigating hallucinations in MLLMs. TARS reformulates direct preference optimization (DPO) as a min-max objective that (1) minimizes behavioral misalignment via preference feedback and (2) maximizes adaptability through perturbations of visual-agnostic tokens. Right: Evaluation on LLaVA-v1.5-13B with preference optimization (PO) (Liu et al., 2023b) and various MLLMs under AMBER benchmark (Wang et al., 2023) shows that TARS surpasses PO baselines and matches GPT-4o (Hurst et al., 2024).
@@ 1 Introduction
Large language models (LLMs) demonstrate strong reasoning capabilities across a broad range of language tasks (Gandhi et al., 2023; Chen et al., 2023; Wang et al., 2025). Building on this foundation, multimodal large language models (MLLMs) integrate visual inputs to enable grounded understanding and vision-language reasoning (Tong et al., 2024; Huang et al., 2023; Driess et al., 2025). Although this integration broadens their applicability in various tasks (Jiang et al., 2024b; Shao et al., 2025), it also introduces key challenges, among which hallucinations are particularly prominent (Kim et al., 2024; Jiang et al., 2024a; Huang et al., 2024). Hallucinations in MLLMs refer to outputs that may appear plausible yet are factually incorrect or lack grounding in the visual context (Sarkar et al., 2025b; Gunjal et al., 2024). Addressing these failures is essential for improving the reliability, safety, and practical applicability of MLLMs in real-world applications.
@@
Modern MLLMs are typically developed through a two-stage training pipeline that includes a knowledge-intensive pretraining phase (Dai et al., 2023; Bao et al., 2022; Zhang et al., 2024), followed by instruction tuning (Liu et al., 2024a; Wang et al., 2024b; Liu et al., 2023a). These stages endow the model with broad world knowledge and the ability to follow instructions in natural language (Bao et al., 2022; Liu et al., 2023b). Despite these capabilities, hallucinations often stem not from knowledge deficits, but from behavioral biases acquired during training that lead the model to generate plausible yet ungrounded outputs (Oh et al., 2024; Chen et al., 2025). To address these failures, preference optimization (PO) has become a prominent strategy for reducing hallucinations by aligning model outputs with human expectations (Schulman et al., 2017; Achiam et al., 2017). PO fine-tunes models using ranked response pairs derived from either human feedback (Sun et al., 2024; Ouyang et al., 2022) or AI-generated preferences (Yu et al., 2025; Sharma et al., 2024), providing direct supervision to reinforce grounded and faithful responses. Such methods have proven effective in mitigating hallucinations across diverse tasks.
Direct preference optimization (DPO) (Rafailov et al., 2023) has become a widely adopted method for hallucination reduction (Fu et al., 2025; Yang et al., 2025). While generally effective, current DPO methods rely heavily on preference data, which can cause models to overfit to shallow textual cues, such as high-frequency phrases or repetitive patterns in the training set (Huang et al., 2024; Liu et al., 2024b). Prior studies have observed that this overfitting leads MLLMs to generate responses that appear plausible but lack visual grounding, as illustrated in Figure 2(c). In our analysis (Figure 2(d)), we further find that DPO-trained models often assign high preference to outputs containing spurious correlation tokens, including prepositions or frequently mentioned objects, even when these elements are not visually grounded (Xie et al., 2024; Wang et al., 2024a). These observations reveal a core limitation of DPO: its reliance on static preference signals hinders generalization under shifting visual–textual contexts, leading to brittle cross-modal alignment and increased vulnerability to hallucination (Setlur et al., 2024; Fu et al., 2025). This rigidity also prevents the model from adapting to local semantic discrepancies, weakening its ability to ground responses in causally relevant visual information.
@@
To validate the effectiveness of our method, we evaluate TARS on LLaVA-v1.5 (Liu et al., 2023b) at 7B and 13B scales, comparing it against leading preference optimization approaches. Across a comprehensive suite of hallucination benchmarks spanning generative and discriminative tasks, TARS-enhanced LLaVA-v1.5 achieves consistently strong performance and matches GPT-4o (Hurst et al., 2024) in several settings. These results underscore the effectiveness of token-adaptive preference optimization in reducing hallucinations and advancing the trustworthiness of MLLMs.
@@ 2 Preliminaries
Multimodal Large Language Models. MLLMs extend LLMs by incorporating visual inputs alongside textual prompts (Zhang et al., 2024). Formally, given an image $x$ and a prompt $q$, the model generates a textual response $y=(y_{1},\dots,y_{l})$ in an autoregressive manner (Liu et al., 2023b):
@@
where $\pi_{\theta}$ denotes the conditional generation policy parameterized by $\theta$. Given a textual input $q$ and a visual input $x$, the model tokenizes them into discrete sequences: textual tokens $q=\{q_{1},\dots,q_{m}\}$ and visual tokens $x=\{x_{1},\dots,x_{n}\}$. These tokens are mapped to embeddings and fused via cross-attention to integrate semantic signals from both modalities. The resulting context is then used by the decoder to autoregressively generate the output sequence (Dou et al., 2022; Yang et al., 2021).
Direct Preference Optimization. Direct preference optimization (DPO) (Rafailov et al., 2023) is an effective approach for aligning model behavior with human preferences. It bypasses explicit reward models by directly optimizing preferences from pairwise comparisons.
Traditional methods such as reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022) and AI feedback (RLAIF) (Yu et al., 2025) rely on training a scalar reward model $r_{\psi}(x,q,y)$ from preference pairs. This reward model is typically trained using the Bradley-Terry formulation (Bradley & Terry, 1952):
@@
$\underset{\pi_{\theta}}{\min}\,\mathbb{E}_{(x,q)\sim\mathcal{D},\,y^{*}\sim\pi_{\theta}(y^{*}\mid x,q)}\Big{[}-\Big{(}r_{\psi}(x,q,y^{*})-\alpha\cdot\mathbb{D}_{\mathrm{KL}}\left(\pi_{\theta}(y^{*}\mid x,q)\,\middle\|\,\pi_{\text{ref}}(y^{*}\mid x,q)\right)\Big{)}\Big{]},$ (4)
where $\alpha$ controls the strength of regularization, which ensures alignment with the learned preferences. Rather than relying on the explicitly trained reward model, DPO (Rafailov et al., 2023) simplifies the learning process by leveraging the insight that the optimal policy can be expressed in closed form using relative log-likelihoods under $\pi_{\theta}$ and $\pi_{\text{ref}}$:
$\underset{\pi_{\theta}}{\min}\,\mathbb{E}_{(x,q,y_{w},y_{r})\sim\mathcal{D}}\Big{[}-\log\sigma\Big{(}\alpha\log\frac{\pi_{\theta}(y_{w}\mid x,q)}{\pi_{\text{ref}}(y_{w}\mid x,q)}-\alpha\log\frac{\pi_{\theta}(y_{r}\mid x,q)}{\pi_{\text{ref}}(y_{r}\mid x,q)}\Big{)}\Big{]},$ (5)
@@ 3 Method
Figure 3: Overview of TARS. TARS reformulates preference optimization as a Min–Max problem: (1) The maximization branch perturbs visual-agnostic tokens to simulate semantically shifted contexts (red dashed box); (2) The minimization branch fine-tunes the model to align with human preferences via the DPO objective (purple dashed box). TARS encourages the model to attend to causally grounded visual signals rather than spurious correlations, thereby reducing hallucinations.
@@ 3.2 Maximizing with Token Perturbations
As shown in Equation 5, DPO aligns models with preferred responses via log-likelihood ratios against a reference model. However, we observe that this formulation can encourage overfitting to superficial patterns such as frequent phrases, stylistic tokens, which we find reduce effective alignment with the visual context in multimodal settings (Setlur et al., 2024; Fu et al., 2025).
@@
where $\Phi(\mathcal{A})$ denotes allowable perturbations constrained to $\mathcal{A}(x,q)$, and $\mathrm{Sim}(\varphi(q),q)$ measures token-level deviation. In practice, we approximate $\varphi^{*}(q)$ by applying token-level transformations:
@@
where $\lfloor\cdot\rfloor$ denotes the floor operation, $\omega$ is a scaling coefficient. $P=-\mathcal{G}_{v}(x)\mathcal{G}_{t}(q_{i})^{T}$ is the similarity score matrix, $\Delta P=\max_{j}P_{j}-\max_{k\neq j}P_{k}$ quantifies the predictive uncertainty of MLLMs. We adapt $N_{t}$ inversely to this margin: confident predictions lead to fewer perturbations, while greater uncertainty induces broader variation.
@@ 3.3 Spectral Regularization for Token Alignment
While our method introduces token-level perturbations to simulate distribution shifts, the supervision derived from preference pairs $(y_{w},y_{r})$ is static. This discrepancy between adaptive input representations and fixed supervision may encourage the model to learn distribution-specific artifacts, especially under strong alignment constraints (Fu et al., 2025; Chowdhury et al., 2024).
In practice, semantic alignment does not require strict token-level correspondence. Enforcing fine-grained constraints may reintroduce spurious correlations that our min–max strategy aims to mitigate (Zhou et al., 2024; Tian et al., 2025). To address this, we propose frequency-domain alignment, where local token perturbations translate into smooth variations in spectral space. This approach ensures semantic consistency between perturbed and inputs without rigid token-wise matching.
Specifically, we extract hidden states for $(x,\varphi(q),y_{w})$ and compare them to those of $(x,q,y_{w})$ and $(x,q,y_{r})$ using the Fast Fourier Transform (FFT) (Cooley & Tukey, 1965). Formally, let $z\in\mathbb{R}^{L\times D}$ denote a sequence of hidden states, where $L$ is the token length. We compute the spectral representation via:
@@
$\mathcal{L}_{\text{freq}}=-\log\sigma\Big{(}\beta\cdot\Big{[}\,\log\frac{\mathcal{F}\left(h_{\theta}(x,\varphi(q),y_{w})\right)}{\mathcal{F}\left(h_{\text{ref}}(x,q,y_{w})\right)}-\log\frac{\mathcal{F}\left(h_{\theta}(x,\varphi(q),y_{w})\right)}{\mathcal{F}\left(h_{\text{ref}}(x,q,y_{r})\right)}\,\Big{]}\Big{)},$ (11)
@@ 3.4 Minimization Objective in TARS
$\mathcal{L}_{\text{TARS}}(x,q,\varphi(q),y_{w},y_{r})=\;\mathcal{L}_{\text{DPO}}(x,\varphi(q),y_{w},y_{r})+\lambda\cdot\mathcal{L}_{\text{freq}}(x,q,\varphi(q),y_{w},y_{r}),$ (12)
@@
4 Experiments
4.1 Experiment Details
Experiment Setups. We evaluate our method using the multimodal large language model LLaVA-v1.5 (Liu et al., 2023b) at both 7B and 13B parameter scales. All evaluations are conducted using greedy decoding with a temperature of 0 to ensure deterministic outputs and reproducibility.
To enable fair comparison, we align our training configuration with the most data-efficient preference optimization baselines. Specifically, we randomly sample 4.8k instances from the RLHF-V-Dataset (Yu et al., 2024), consistent with OPA-DPO (Yang et al., 2025), and adopt the same training strategy as CHiP-DPO (Fu et al., 2025). All models are trained on eight NVIDIA A100 (80GB) GPUs. We set $\alpha=1$ in Equation 5 and $\beta=1$ in Equation 11 for preference optimization. We implement $\varphi(\cdot)$ using both token masking (Mask) and replacement (Replace) strategies in Equation 8, and set the perturbation constraint strength to $\omega=0.1$ in the adversarial min-max formulation Equation 9. We use a frequency-domain loss weight of $\lambda=0.1$ in Equation 12. Full implementation details and ablation studies are provided in the Appendix.
Evaluation Benchmark. We evaluate TARS across both generative and discriminative hallucination benchmarks to ensure that hallucination mitigation does not come at the cost of factual grounding. Our evaluation framework includes four established benchmarks:
1) AMBER Wang et al. (2023) (Generative): A comprehensive 1004-image benchmark targeting fine-grained evaluation of hallucinations. In line with prior works Yang et al. (2025); Fu et al. (2025), we evaluate only the generative subset using the official codebase. Metrics include CHAIR Rohrbach et al. (2018), object coverage (Cover), response-level hallucination rate (Hal-Rate), and alignment with human cognition (Cog).
2) MMHal (Sun et al., 2023) (Generative): A VQA benchmark with 96 real-world scenarios, evaluated using GPT-4V feedback to measure overall scores and hallucination rates (Hal-Rate).
3) OBJHal (Yu et al., 2024) (Generative): A 300-instance benchmark evaluating hallucinations in captioning. We report hallucination rates at the response level (CRs) and object mention level (CRi).
4) POPE (Li et al., 2023) (Discriminative): A 9000-sample binary VQA benchmark designed to assess object hallucination along the textual axis through yes/no questions.
Baseline Methods. We compare against two categories:
(1) Advanced multimodal foundation models: Intern-VL2.5-7B (Chen et al., 2024e), Qwen-VL2.5-8B (Bai et al., 2025), DeepSeek-VL2-27B (Wu et al., 2024), and GPT-4o (Hurst et al., 2024) serve as reference benchmarks.
(2) LLaVA-v1.5 with RL techniques: We evaluate multiple RL-based approaches applied to both the 7B and 13B variants of LLaVA-v1.5, including RLHF (Sun et al., 2024), RLAIF (Yu et al., 2025), HALVA (Sarkar et al., 2025a), as well as three state-of-the-art methods based on direct preference optimization (DPO): DPO (Pi et al., 2024), CHiP-DPO (Fu et al., 2025), and OPA-DPO (Yang et al., 2025). A comparison of algorithmic properties is provided in Table 3.
Table 1: Comparison across hallucination benchmarks. We evaluate SOTA MLLMs as reference baselines, denoted by $\S$. For algorithms with available checkpoints, results from re-testing are marked with $\dagger$; for those without, we reproduce results using settings from (Fu et al., 2025; Li et al., 2024), denoted by $\ddagger$. All experiments use greedy decoding with temperature set to 0 for consistency and reproducibility. Bold denotes the best performance, and underlined denotes the second-best.